OpenStudy (annie96):

Solve, using the substitution method. j + k = 3 j – k = 7 (Points : 1) The solution is (8, 1) The solution is (5, –2) There is no solution. There are an infinite number of solutions.

jigglypuff314 (jigglypuff314):

j + k = 3 j – k = 7 --> j = k + 7 j + k = 3 --> (k + 7) + k = 3 k + 7 + k = 3 2k + 7 = 3 -7 -7 2k = -4 k = -2 <-- (which option has a -2 in it? :P )

OpenStudy (anonymous):

the solution is (5,-2)

OpenStudy (annie96):

Thank you guys.